### v1

# Revisiting the question of what instantaneous normal modes tell us about liquid dynamics

Abstract
The absence of a well-defined equilibrium reference configuration and the inevitable frequent atomic rearrangements have long obstructed the achievement of a complete atomic-level understanding of liquid dynamics and properties, a challenge that continues to be unresolved. The instantaneous normal mode (INM) approach, based on the diagonalization of the potential energy Hessian matrix in instantaneous liquid configurations, has been shown to be a promising starting point to predict thermodynamic and dynamical properties of liquids but presents several conceptual difficulties that remain to be addressed. More in general, due to the inability of capturing anharmonic effects, what INMs can tell us about liquid dynamics remains an open question. In this work, we provide a general set of “experimental facts” about the INM spectrum of liquids by performing a comprehensive INM analysis of several simulated systems, including Ar, Xe, N2, CS2, Ga and Pb, in a large range of temperatures from the solid to the gas phase. We first study the INM density of states (DOS) and compare it to the density of state function obtained from the velocity auto-correlation function. Secondly, we analyze the temperature dependence of the fraction of unstable modes and of the low-frequency slope of the INM DOS, in search of possible universal behaviors. We then explore the connection between INMs and other dynamical and structural properties of liquids including the liquid-like to gas-like dynamical crossover and the momentum gap of collective shear waves. Moreover, we investigate the INM spectrum at low temperature and its properties upon approaching the solid phase, revealing the existence of a large fraction of unstable modes also in crystalline solids. Finally, we verify the existence of a recently discussed cusp-like singularity in the INM eigenvalue spectrum and reveal its complex behavior upon dialing temperature that challenges the existing theoretical models.
preprint: AIP/123-QED
I Introduction
Normal modes describe the coherent motion of atoms oscillating at the same frequency and they represent the basis for the study of vibrations in many physical systems from buildings and bridges to molecules and solid state materials Cui and Bahar (2005); Dykeman and Sankey (2010). The most general dynamics in the linear regime can be indeed reconstructed using a superposition of normal modes that act as building blocks for thermodynamic and transport properties as well. A paradigmatic example is that of harmonic crystals in which the normal mode analysis provides a complete atomic-scale description Moon (2023), and leads to the well-established and successful Debye Debye (1912) and Einstein Einstein (1907) models.
An analogous microscopic description of the physical properties of liquids incurs in several difficulties (see Moon for an extensive discussion) and remains unachieved Cabrillo et al. (2002). Consistent with Maxwell’s idea Maxwell (1867) that liquids behave as solids below a certain time-scale (Maxwell relaxation time), in 1967 Zwanzig proposed that normal modes could also exist in liquids, and be a fundamental part of the dynamics Zwanzig (1967). This hypothesis was quickly verified by Rahman, Mandell, and McTague Rahman et al. (1976) using a LJ simulated model, giving birth to the concept of “instantaneous normal modes” (INM), that was developed later on to great extent by many Authors Keyes (1997a).
INM can be obtained by probing the potential energy surface (PES) at any instant of time and diagonalizing the corresponding instantaneous Hessian matrix that encodes all the information about the local curvature of the PES. In this way, the INM density of states (DOS) can be defined by averaging on several of these instantaneous configurations. Despite nowadays this procedure can be easily achieved using computer simulations, what INM are and are not remains a challenging question Stratt (1995). In particular, whether the harmonic oscillator picture is just a mere mathematical formal analogy or some kind of coherent oscillatory motion occurs in liquids is a topic of debate. The INM analysis assumes that interactions are harmonic (fully captured by the second derivative of the potential – Hessian matrix), which is a very questionable assumption in liquids. Nevertheless, this crude approximation might have some validity if the Maxwell picture, where short time dynamics is elastic in nature, is accepted. Importantly, short time dynamics has repeatedly been found to be directly linked to long time dynamics so that the limited applicability of the INM analysis might conceivably be extended to longer time relaxation processes. [^1] We would like to thank Jack F. Douglas for suggesting this point to us.
Among the various difficulties in this program, it was early realized by Rahman Mandell, and McTague Rahman et al. (1976) that, despite the INM density of states displays striking similarities with the Fourier transform of the velocity auto-correlation function (VACF), the INM procedure inevitably leads to the presence of negative eigenvalues corresponding to regions in the PES with locally negative curvature. These negative eigenvalues translate into imaginary frequencies, i.e., “unstable modes”. INM can be therefore divided into stable (positive eigenvalues and real frequencies) and unstable (negative eigenvalues and imaginary frequencies), and the same separation can be applied to the corresponding INM DOS. Let us immediately remark that the “density of states” of unstable INM cannot be directly interpreted as a vibrational density of states. More in general, unstable modes cannot be identified with collective atomic vibrations and their physical interpretation appears more subtle and related to the relaxational processes, active in liquids but suppressed in crystals.
The fact that liquids display solid-like oscillatory dynamics at large frequencies, or equivalently short times, it is not a mystery nor a surprise Vallauri and Bermejo (1995), and it is experimentally verified, e.g., Giordano and Monaco (2010). These fast dynamics can be well captured by the INM analysis and by the concept of stable INM, whose density of states is in good agreement with the Fourier transform of the VACF above a certain cutoff frequency (see for example Seeley et al. (1992)). On the other hand, it is well-know that these two quantities are different at low frequencies (compared to an inverse characteristic time $1/\tau$). As a concrete example, the zero frequency value of the INM DOS vanishes while the corresponding limit in the VACF gives the non-zero self-diffusion constant. This is just a manifestation that a harmonic description of liquids cannot apply at long times because of the inevitably anharmonic nature of the potential landscape David and Stratt (1998) (some anharmonic extensions of the INM approach have been proposed in the past, e.g., Krämer et al. (1998)). It was therefore a surprising discovery Keyes (1994a, 1995) that INM, and in particular unstable ones, can still carry important information about the late-time liquid dynamics and that the self-diffusion constant can be reconstructed from the distribution of some of them (even if a concrete characterization of the adjective “some” is a challenging task, see Gezelter et al. (1997) and related discussions Li et al. (1998); Chowdhary and Keyes (2002); Gezelter et al. (1999); Clapa et al. (2012a); Madan et al. (1990a)). More recently, it has been proposed that another transport property of liquids – the shear viscosity – can be reconstructed from unstable INMs Huang et al. (2024).
In frequency domain, the INM density of states exhibits interesting features that have been subject of intense investigation in the past decades. Particular attention has been devoted to understanding the frequency and temperature dependence of the unstable INM DOS Keyes et al. (1997); Vijayadamodar and Nitzan (1995); Zürcher and Keyes (1997); Xu et al. (2023), which is commonly represented on the negative frequency axes upon performing a variable transformation $\omega=-\sqrt{|\lambda|}$ with $\lambda$ the (negative) eigenvalues. Among the various features, the unstable INM DOS displays a universal linear in frequency scaling at low frequency that is observed in uncountable simulations and derived within several theoretical frameworks Xu and Stratt (1989); Wu and Loring (1993); Wan and Stratt (1994); Zaccone and Baggioli (2021); Schirmacher et al. (2022); Li and Keyes (1999); Zürcher and Keyes (1997); Wu and Loring (1992). It is nowadays clear that the emergence of this scaling is intimately connected to the existence of unstable modes. Moreover, by construction the stable and unstable INM DOS are symmetric upon the transformation $\omega\rightarrow-\omega$, implying that also the stable branch of the INM DOS displays a linear scaling with the same temperature dependent coefficient. Additional interesting properties emerge when the INM DOS is represented in function of the eigenvalues, rather than frequencies. In particular, it was early observed Sastry et al. (2001); Taraskin and Elliott (2002), and recently re-considered Schirmacher et al. (2022); Mossa et al. (2023), that the eigenvalue distribution presents a cusp-like singularity at $\lambda=0$ that remains hidden in the frequency representation but carries important information about liquid dynamics as well.
The relation (if any) between the INM DOS, the Fourier transform of the VACF (sometimes called the “density of state function”) and the experimentally measurable density of states of liquids remains an important open question. The density of states of liquids at low energies is much less explored experimentally compared to the solid counterpart as the standard experimental probes incur in several difficulties. Until very recently, only few works explored the low-frequency DOS of liquids using inelastic neutron scattering techniques, i.e., Ref. Phillips et al. (1989) for liquid selenium and Ref. Dawidowski et al. (2000) for heavy water. More recently, the low-frequency experimental DOS of different liquid systems was investigated in Ref. Stamper et al. (2022) and a universal linear scaling was observed in agreement with the expectations from the INM analysis. Furthermore, the temperature dependence of the linear scaling of the experimental DOS was studied in Ref. Jin et al. (2024) and compared to the linear scaling of the INM DOS. Both scaling coefficients were found to follow the same exponential dependence on the inverse temperature; nevertheless, the corresponding activation energies were found to be different. Despite these encouraging developments, how to define “a” liquid density of states, and if such a quantity ultimately exists, remain works in progress.
Arguably, the difficulties in performing a normal mode analysis for liquids, and in identifying a well-defined and unique density of states, are reflected in the challenge of achieving a microscopic description of liquid thermodynamics and collective dynamics Trachenko and Brazhkin (2015a). In the last years, several works Baggioli and Zaccone (2021); Xu et al. (2023); Moon et al. (2024, 2023) have attempted to use INMs to describe the thermodynamic properties of liquids and in particular their heat capacity. From these analyses, it has emerged that unstable modes cannot be ignored in liquids as considering only stable normal modes would lead to erroneous conclusions. This is somehow consistent with the older two-phase description of liquids, e.g. Lin et al. (2003), when the gas-like modes are related to the unstable INM modes as done in Ref. Moon et al. (2024). Interestingly, unstable INM have started to play an important role in glasses and amorphous solids as well Kriuchevskyi et al. (2022); Oyama et al. (2021). For example, unstable INM can be used to locate the onset of glass transition Bembenek and Laird (1995). Finally, it is not yet clear how the INM picture developed for liquids merges into the more traditional normal mode description of solids, for which Debye’s law has to be recovered. It is indeed known Pallikara et al. (2022), but often discarded, that also crystalline systems at finite temperature possess a non negligible number of unstable modes.
In summary, there is increasing evidence that instantaneous normal modes might provide a basis for a microscopic description of liquid properties from their atomic-level motion, to their thermodynamics and collective properties. Nevertheless, the question of what we can learn from INM (in particular the unstable ones) is still open, see e.g. Zhang et al. (2019). In this work, we revisit this question and we provide a comprehensive INM analysis for many liquid systems in order to collect a series of “experimental facts” (hopefully) useful to build the long-sought microscopic theory of liquids.
In Section II, we present our simulation models and the INM analyses carried out. In Section III, we take Argon as a case study and we perform an extensive analysis using several methods to explore the role of INM in such a system. We then extend the analysis to a large class of liquids including Xe, N2, CS2, Ga and Pb in order to verify the universality of our findings. In Section IV, we select a set of topics that are part of the broader question of what we can learn from INM about liquids and we provide several “experimental facts” within this debate. Finally, in Section V, we conclude and discuss the most interesting open directions following from our results.
II Methods and models
We start by describing in full detail the simulations models considered in our study and the methods used.
II.1 Interaction potential
In this work, all simulations are conducted via the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Plimpton (1995). The computer simulation of metals was conducted using the pair interparticle interaction method, along with the incorporation of a more advanced embedded atom model (EAM) potential technique. In the EAM, an additional form of collective interaction is introduced alongside the conventional pair interparticle potentials. The expression for the potential energy of a metal is given as
$U=\sum_{i}\Phi(\rho_{i})+\sum_{i<j}{\phi(r_{ij})},$ (1)
where $\Phi$$(\rho_{i})$ is the potential of embedding the $i$th atom depending on the "effective electron density" $\rho_{i}$ at the location of the atom’s center and it is determined by the surrounding atoms
$\rho_{i}=\sum_{j}\Psi(r_{ij}),$ (2)
where $\Psi$$(r_{ij})$ is the contribution to electron density of neighbor atom $j$. The second term in Eq. (1) represents the usual pair potential. In this work, the interactions between metal crystals of Ga Belashchenko (2012) and Pd Wang et al. (2018) atoms are described with EAM potentials independently, which are widely applied in respective metallic systems Mokshin et al. (2015); Chtchelkatchev et al. (2016); Belashchenko (2017); Wonglakhon et al. (2021). The rest of non-metal systems of particles interact via the 12-6 Lennard-Jones (LJ) potential,
$\phi_{\alpha\beta}(r_{ij})=4\epsilon_{\alpha\beta}\left[\left(\frac{\sigma_{\alpha\beta}}{r_{ij}}\right)^{12}-\left(\frac{\sigma_{\alpha\beta}}{r_{ij}}\right)^{6}\right],$ (3)
where $i$ and $j$ stand for $i$th and $j$th atom, $\alpha$ and $\beta$ for the type of atoms, and $\epsilon_{\alpha}$β and $\sigma_{\alpha}$β are the LJ parameters. All parameters chosen to approximate the intermolecular potential for each system are list in Table 1. For nitrogen molecules, the intramolecular interaction is approximated by the harmonic type potential as
$U_{NN}(r)=\frac{1}{2}K_{NN}(r-r_{e}),$ (4)
where $K_{NN}$ is the force constant which is measure of interatomic bond between nitrogen atoms in N2 molecule, $r$ is the bond length, and $r_{e}$ is the bond length at equilibrium. We used a bond spring constant $K_{NN}$ = 13800 KJ/(mol) and a bond length $r_{e}=1.0975$. At the same time, the harmonic type potential for CS2 molecules S.-B. Zhu and Robinson (1988) is given by
$U_{intra}=\sum_{bonds}\left[K_{1}\{\left(r_{1}-r_{0}\right)^{2}+\left(r_{2}-r_{0}\right)^{2}\}+K_{2}\left(r_{1}-r_{0}\right)\left(r_{2}-r_{0}\right)\right]+\sum_{angles}K_{3}\left(\theta-\theta_{0}^{2}\right)^{2},$ (5)
where $r_{1}$ and $r_{2}$ are the bond lengths of C-S, and $\theta$ is the bond angle of S-C-S in the CS2 molecule. The parameters used in equation (5) are shown in Table 2.
Table 1: Lennard-Jones energy and length parameters.
[…]
Ar Mourits and Rummens (1977) ArAr 3.465 0.225 8
Xe Mourits and Rummens (1977) XeXe 4.05 0.457 10
N2 Gao and Weiner (1990) NN 3.31 0.0715 12
CS2 Tildesley and Madden (1981) CC 3.35 0.1017
SS 3.52 0.3637 12
CS 3.44 0.1923
Table 2: The intramolecular potential parameters for the CS2 model.
Parameters Values
$K_{1}$ 2250 KJ/(molÅ$\cdot\textup{\AA}^{2}$)
$K_{2}$ 360 KJ/(molÅ$\cdot\textup{\AA}^{2}$)
$K_{3}$ 169 KJ/(molÅ$\cdot\textup{\AA}^{2}$)
$r_{0}$ 1.55 Å
$\theta_{0}$ 180
II.2 Simulation details
For the four monatomic systems, a FCC lattice structure of 4000 atoms with lattice constant of $a_{A}$r=$5.268\,\textup{\AA}$ Henshaw (1958), $a_{X}$e= $6.347\,\textup{\AA}$ Granfors et al. (1981), $a_{G}$a= $4.4\,\textup{\AA}$ Belashchenko (2012) and $a_{P}$b= $4.95\,\textup{\AA}$ Wang et al. (2018) was respectively used as an initial input structure for the MD simulations. For N2 and CS2 systems, the MD simulations were performed with 500 N2 and 3375 CS2 molecules in either one cubic box of dimension 30.0 Å and 80.0 Å respectively. All samples were simulated with periodic boundary conditions (PBC). In the process of samples preparation, after energy minimization was performed using steepest descent algorithm to bring the initial configuration in minimum potential, all samples relaxed at least $5$ ns (MD time step is $1.0$ fs) to reach their respective equilibrium at a series of temperatures and specific pressures based on phase diagram to keep them liquid. In this process, the isobaric and isothermal (NPT) ensemble was used with the sample size being adjusted to give target pressure. After that, the canonical (NVT) ensemble MD was conducted at different temperatures for data collections and analysis with an interval of $100$ fs during a $1$ ns simulation. The MD simulation systems of Ar and CS2 are shown in Figure 1.
Figure 1: MD simulation systems for liquid Ar and CS2. The snapshot of liquid Ar (a) at $T=100$ K, $P=40$ bar and liquid CS2 (b) at $T=300$ K, $P=1.01$ bar. Figures made with OVITO software Stukowski (2009).
II.3 Definition of the physical observables
In this Section, precise definitions of the key physical quantities considered in this work are provided.
The velocity auto-correlation function (VACF), $Z(t)$, is an essential time-correlation function in physics and chemistry that provides extensive information about molecular-structural and hydrodynamic properties of matter Hansen and McDonald (2013). The Fourier transform of $Z(t)$ is defined as the density of state function, $S(\omega)$. More precisely,
[…]
On the other hand, the instantaneous normal mode analysis provides an instantaneous picture of liquid dynamics that has been widely used in the past Keyes (1997a). The INM method involves calculating the normal modes of a system at a given instantaneous configuration, providing insight into the vibrational properties and stability of the system at specific moments in time. It starts from the dynamical matrix that characterizes the curvatures of the local potential energy surface. The dynamical matrix is a $3N\times 3N$ matrix whose elements are defined as follows:
$H_{i\mu,j\nu}(\mathbf{R})=\frac{1}{\sqrt{m_{i}m_{j}}}\frac{\partial^{2}V}{\partial r_{i,\mu}\partial r_{j,\nu}}.$ (7)
Here, $i,j=1,...,N,\mu,\nu=x,y,z$. $N$ is the number of total atoms in the system. $\mathbf{R}\equiv{\mathbf{r}_{1},...,\mathbf{r}_{N}}$ represents each liquid configuration and $\mathbf{r}_{i}$ is the position of the $i$th atom. V is the potential energy and $r_{i,\mu}$ represents the $\mu$-coordinate of the $i$th atom. The instantaneous normal mode frequencies $\omega_{i}$ are defined as the square roots of the eigenvalues of the dynamical matrix $\lambda_{i}$. The INM density of states, $\rho(\omega)$, is then obtained by averaging over different instantaneous configurations,
$\rho(\omega)=\left<\frac{1}{3N}\sum_{i}^{3N}\delta(\omega_{i}-\omega)\right>.$ (8)
Due to locally negative curvature regions in the potential energy landscape, INMs can be divided into stable modes (positive eigenvalues) and unstable modes (negative eigenvalues). We define the fraction of unstable modes, $f_{u}$, as follows:
$f_{u}=\frac{N_{u}}{3N},$ (9)
where $N_{u}$ presents the number of unstable modes. The complementary fraction of stable mode $f_{s}$ is just its complement, i.e. $f_{s}=1-f_{u}$. In an ideal zero temperature crystal, one expects $f_{u}=0$ and $f_{s}=1$.
Finally, we define the slope of the INM DOS as $a(T)$ from the low-frequency behavior,
$\rho(\omega)=a(T)|\omega|+\dots$ (10)
where negative values of $\omega$ correspond to imaginary frequencies, and the ellipsis stands for higher order corrections in $\omega$.
From the MD simulation, we also calculate the diffusion coefficient, $D$, and the pair correlation function, $g(r)$, to gain a comprehensive understanding of the dynamic and structural properties of the simulated systems. The diffusion coefficient, $D$, quantifies the rate of particle diffusion and is determined using the Einstein relation, which relates $D$ to the particle mean square displacement (MSD),
$D=\lim_{t\to\infty}\frac{1}{6t}\left<\lvert\mathbf{r}(t)-\mathbf{r}(0)\rvert^{2}\right>,$ (11)
[…]
$g(r)=(V/N)\left[n(r)/4\pi r^{2}\Delta r\right]$ (12)
Here, $V$ represents the volume of the system, $N$ is the total number of particles, and $n(r)$ is the number of atoms that are situated at a distance between $r$ and $r+\Delta r$ from a given particle. The term $4\pi r^{2}\Delta r$ is the volume of the spherical shell at distance r with thickness $\Delta r$. Physically, $g(r)$ reveals how the probability of finding a particle at a distance $r$ from a reference particle deviates from the ideal gas limit. Precisely, $g(r)=1$ indicates that the particles are distributed randomly at distance $r$, while $g(r)$ greater than $1$ suggests the presence of local structural order. Conversely, $g(r)$ less than $1$ indicates a depletion of particles at that distance, suggesting the presence of repulsive interactions or an exclusion zone around each particle.
III Argon as a benchmark case
We take the case of Argon (Ar) as a paradigmatic simple example to study liquid properties within the INM framework. The phase diagram of Argon in the regime of interest is shown in Fig.2. The black line indicates the line of first-order phase transitions separating the liquid (light blue) and the gas (light green) phases. The critical point, indicated with a red circle, is located at $T_{c}\approx 151$ K and $P_{c}\approx 48.55$ bar Mccain and Ziegler (1967). The data for the phase diagram are collected from the National Institute of Standards and Technology (NIST) database Standards (2005). The pink region is the supercritical phase beyond the critical point. The horizontal red isobar lines indicate some of the regions that have been numerically investigated in this work. Other investigated isobar lines with much higher pressure are not displayed in Fig.2. Finally, the blue circles indicate the onset of the dynamical crossover (Frenkel line) that will be discussed in detail below.
Because of space limitations, in the main text we focus only on the isobar line at $P=40$ bar and we explore a wide region of temperatures from $95$ K to $160$ K. This isobar line crosses the first order phase transition between liquid and gas at approximately $T\approx 146$ K.
Figure 2: A zoom of the phase diagram of Argon (Ar) indicating the liquid and gas phases as well as the critical point at approximately $151$ K and $48.55$ bar. The horizontal red lines indicate the set of thermodynamic conditions explored in this manuscript. The blue dots correspond to the dynamical crossover points discussed below.
We start our analysis by considering the behavior of the velocity auto-correlation function (VACF) $Z(t)$ and the pair distribution function (PDF) $g(r)$ that are shown in panels a and b of Fig.3 respectively. At high temperature, deep in the gas regime, $Z(t)$ decays monotonically and its form is well approximated by an exponential function $\exp(-\gamma t)$ (shown with black dashed line), as predicted by kinetic theory under the assumption of pure Brownian motion Hansen and McDonald (2013). Upon decreasing the temperature, $Z(t)$ remains a monotonically decreasing function but it is not anymore well approximated by a pure exponential function. This indicates the onset of correlations and the tendency towards dense liquid dynamics. A big jump is observed between $140$ K and $130$ K that corresponds approximately to the boiling temperature of Argon at this pressure. Even in the liquid phase, at least up to a certain “critical” temperature, the velocity auto-correlation function remains a monotonically decreasing function.
More importantly, as emphasized in the inset of panel a in Fig.3, the velocity-autocorrelation function displays a drastic change of behavior around $115$ K where it starts developing a minimum that transforms into marked oscillations by decreasing the temperature further. This drastic change of behavior corresponds to a transition between gas-like dynamics to liquid-like dynamics and it is known in the literature as the Frenkel line Brazhkin and Trachenko (2012); Brazhkin et al. (2012, 2013). The Frenkel line has been studied in several liquids in the supercritical phase Yoon et al. (2018); Bolmatov et al. (2013a, 2014); Prescher et al. (2017); Bolmatov et al. (2015); Fomin et al. (2018); Proctor et al. (2019); Fomin et al. (2015); Brazhkin et al. (2011); Cockrell et al. (2021), but also in complex dusty plasmas Huang et al. (2023); Ge et al. (2023) and active granular fluids Jiang et al. (2024). Here, we report the existence of this dynamical crossover also below the critical point, away from the supercritical region. This confirms, as already noted in the previous literature, that the Frenkel line does not emanate from the critical point and it persists also in the liquid phase. We have tracked the point at which this dynamical crossover appears for three different pressures. The results are indicated with blue symbols in the phase diagram, Fig.2. From there, it is clear that the Frenkel line extends into the liquid phase and goes almost parallel to the line of first-order phase transitions below the critical point.
Figure 3: (a) The velocity auto-correlation function (VACF) $Z(t)$ defined in Eq.(6) for Argon at $P=40$ bar and different values of temperature. The inset zooms on the temperature values below $T=140$ K to emphasize the dynamical crossover (Frenkel line) that happens at $T\approx 115$ K. The dashed black lines are fits to a purely exponential function as discussed in the main text. (b) The corresponding pair correlation function $g(r)$. The arrow indicates the structural signatures in the second peak of $g(r)$ of the first order gas-liquid phase transition that happens between $130$ K and $140$ K. The inset shows the disappearance of intermediate range order (third peak in $g(r)$), whose onset is consistent with the dynamical crossover in the inset of panel (a).
In order to investigate structural properties, in panel b of Fig.3 we have computed the pair distribution function $g(r)$ in the same range of temperatures and constant pressure $P=40$ bar. At low temperatures, we observe the typical behavior of a dense liquid with a first strong peak around $4\,\textup{\AA}$ and other weaker peaks around $7\,\textup{\AA}$, $10.5\,\textup{\AA}$ and so on. By increasing the distance $r$, the peaks disappear signaling the absence of long-range order, as one might expect in a liquid. As shown in the inset of panel b of Fig.3, by increasing the temperature but still below the boiling temperature of $\approx 140$ K, we observe that the higher peaks (above the first two) corresponding to intermediate-range order disappear. For example, as indicated by the black arrows, the intensity of the third peak located at approximately $r\approx 10.5\,\textup{\AA}$ vanishes around $115-120$ K. Interestingly, these temperatures are extremely close to the location at which oscillations in the VACF $Z(t)$ disappear (panel a in Fig.3). As already proposed in Bolmatov et al. (2013b), this suggests the existence of a structural change in the intermediate-range order that might be associated to the Frenkel line.
By increasing further the temperature, across the boiling transition, the structure of the system changes drastically. First, all the peaks apart from the first one disappear, implying that intermediate-range order is totally lost. Second, the minimum around $r\approx 5\,\textup{\AA}$ disappears as well and $g(r)$ displays a sharp jump at that length scale. Finally, despite the height of the first peak is reduced, it does not vanish. This implies that we are still far from the totally uncorrelated ideal hard-sphere gas regime.
After this initial exploration, we move to study the instantaneous normal modes in the same temperature regime with the idea of comparing the INM density of states with the Fourier transform of the VACF (density of state function), denoted as $S(\omega)$. In this analysis, we have extended the range of temperatures below the melting temperature to explore the solid phase near melting as well. At the same time, we have stopped the analysis at $120$ K without going deep into the gas phase where the differences between the two quantities are too large. All the results in the solid phase are kept in black color to avoid confusion. The color scheme for the other temperatures is kept the same as in Fig.3. In panel (a) of Fig.4, we show the density of INM using the standard convention of presenting unstable purely imaginary frequencies on the negative frequency axes. Starting from the highest temperature analyzed, $T=120$ K, we can already identify several interesting features. For all temperatures in the liquid phase, both the stable and unstable branches of the INM density of states are linear in frequency at low frequency and they develop a maximum at intermediate frequencies. For the unstable branch, the maximum is located at approximately $2.5$ ps and its location is almost independent of temperature in the whole liquid phase despite its intensity decreases by decreasing temperature. For the stable branch, the location of the maximum moves slightly to higher energies by decreasing the temperature. Below the melting temperature, we observe still a non-vanishing number of unstable modes but a drastic jump in their density. Moreover, the peak in the stable branch jumps quickly to higher energies and approaches an approximately constant value of $5$ ps.
Figure 4: (a) The INM DOS of Argon at $P=40$ bar and different temperatures as defined in Eq.(8). The stable branch is plotted on the positive frequency axes, while the unstable branch on the negative axes using the standard notation described in the main text. The first-order melting transition happens between $92$ K and $95$ K. (b) The density of state function $S(\omega)$ defined in Eq.(6) from the Fourier transform of the VACF for the same system. Black lines correspond to the solid phase.
In panel b of Fig.4 we show the results for the density of state function $S(\omega)$ defined as the Fourier transform of the velocity auto-correlation function $Z(t)$, Eq.(6). In the solid phase (black lines below $95$ K), the density of state function vanishes at zero frequency, as expected for a solid with zero self-diffusion constant. The low-frequency behavior of $S(\omega)$ in the solid phase is compatible with Debye’s law and shows a characteristic quadratic scaling $S(\omega)\propto\omega^{2}$. Moreover, the prefactor in this scaling law grows by increasing the temperature. This can also be rationalized using the Debye model since that prefactor is inversely proportional to the cube of the average speed of sound $\bar{v}^{3}$, and the latter clearly decreases by increasing temperature because of the softening of the material. Moving away from the low frequency regime, $S(\omega)$ displays a peak in the solid phase that is located approximately at $\omega=5$ ps-1 and it is rather insensitive to temperature. A second shoulder is visible around $\omega=9$ ps-1, followed by a monotonic decay.
At approximately $95$ K, the system enters the liquid phase and the behavior of the density of state function $S(\omega)$ changes drastically. First, the zero frequency value of $S(\omega)$ is not anymore zero as a sign of a finite self-diffusion constant that grows with temperature. Moreover, $S(\omega)$ is not anymore monotonic. It first decays monotonically up to a minimum located around $\omega=1$ ps-1, then, it grows up to a peak at approximately $2.5$ ps-1 and it then decays monotonically again. This behavior is typical of dense liquids, see for example Fig.1 in Ref. Lin et al. (2003). By increasing further the temperature, both the minimum and the peak at low frequency disappear and the density of state function becomes a monotonically decaying function as expected for dilute liquids and gases. By increasing temperature, $S(\omega)$ gets closer to the Lorentzian function predicted by kinetic theory, a sign of the complete lack of correlations.
We can already anticipate some clear differences and similarities between the stable INM density of states, $\rho(\omega)$ with $\omega>0$, and the density of state function $S(\omega)$. For brevity, we will indicate the positive ($\omega>0$) branch of $\rho(\omega)$ as $\rho_{s}(\omega)$ in the rest of the manuscript, where the subscript stands for “stable”. Let us start from the solid phase near melting (black lines). Even there, $\rho_{s}(\omega)$ and $S(\omega)$ are profoundly different below $\approx 4$ ps-1. In particular, while $\rho_{s}(\omega)$ is linear in frequency, $S(\omega)$ is quadratic, as Debye’s model predicts for solids. The origin of this difference can be found in the existence of a large number of unstable modes, that also persists in the heated crystal away from the zero temperature harmonic limit. These remaining unstable modes possibly correspond to local anharmonicities that remain present in crystals at finite temperature. Above a certain critical frequency, $\rho_{s}(\omega)$ and $S(\omega)$ become more similar, showing the same qualitative trend. We notice that the positions of the peak in $\rho_{s}(\omega)$ and $S(\omega)$ are compatible in the solid phase. We also emphasize that the fact that the INM analysis does not capture correctly the long-time dynamics is not a surprise and it is a direct consequence of the instantaneous approximation adopted. The differences between $\rho_{s}(\omega)$ and $S(\omega)$ are expected to vanish by decreasing the temperature or increasing the pressure, as discussed in detail in Seeley et al. (1992). Below, we will show this trend concretely for another liquid system, Gallium.
By increasing temperature into the liquid phase, even qualitatively, $\rho_{s}(\omega)$ and $S(\omega)$ are totally different. In particular, $S(\omega)$ acquires a finite zero frequency value, corresponding to the onset of self-diffusion, while $\rho_{s}(0)$ remains zero. Moreover, $S(\omega)$ becomes non-monotonic, or even monotonically decaying with frequency, while $\rho_{s}(\omega)$ remains linear in frequency at small frequency.
Figure 5: (a) The fraction of unstable modes $f_{u}(T)$ and the self-diffusion constant $D(T)$ as a function of temperature for Argon at $P=40$ bar. The melting temperature is between $92$ K and $95$ K (dashed red line) and the Frenkel line is located approximately around $115$ K (dashed blue line). (b) The slope of the low-frequency linear scaling of the INM DOS $a(T)$ for the same system and conditions.
We now move to discuss in detail the features of the INM density of states. We are mainly interested in the fraction of unstable modes $f_{u}(T)$ defined in Eq.(9) and the coefficient of the low-frequency linear scaling $a(T)$, as determined by fitting the INM DOS using Eq.(10). The fraction of unstable modes is shown as a function of temperature in panel a of Fig.5. $f_{u}$ grows with temperature, as one might expect for a parameter that quantifies the “fluidity” of a system. As already anticipated, $f_{u}$ is not zero in the solid phase near melting but accounts for approximately $10\%$ of the INM. $f_{u}$ displays a sharp peak at the melting temperature between $92$ K and $95$ K, where it changes abruptly from $\approx 12\%$ to $\approx 28\%$. This type of behavior was already observed in the past literature Gezelter et al. (1997). Above melting, in the liquid phase, the fraction of unstable modes grows with temperature following an approximately linear trend and it does not show any feature at the dynamical crossover (Frenkel line) reported in panel a of Fig.3. In panel a of Fig.5, we show the self-diffusion constant $D$ as well. Below the melting transition (dashed red line), but not too far from it, the self-diffusion constant becomes very small, as one might expect in crystals Mehrer (2007). This implies that most of the corresponding non-vanishing unstable modes should not be of diffusive type. In the past, these remaining unstable modes have been associated to so-called “false barriers” Gezelter et al. (1997) or simply described as localized unstable modes Bembenek and Laird (1995); Clapa et al. (2012b). Above the melting temperature, the diffusion constant becomes large and grows with temperature monotonically. Within the precision of our data, we do not observe any clear change of behavior across the Frenkel temperature around $115$ K (blue dashed line), despite $D$ seems to display a quite sharp jump around that temperature. A more in detail investigation is needed to further clarify this point.
In panel b of Fig.5, we show the temperature dependence of the linear scaling of the INM DOS, $a(T)$. This scaling coefficient is the same for the stable and unstable parts of the INM spectrum. $a(T)$ grows monotonically with temperature and displays a sharp peak around the melting temperature. Interestingly, its trend between $95$ K and $100$ K seems to follow the behavior of the self-diffusion constant. In particular, in that range of temperatures both $D$ and $a(T)$ are approximately constant before approaching a more steady growth. On the other hand, we do not observe any feature in $a(T)$ upon crossing the Frenkel crossover.
Figure 6: (a) The momentum gap ($k$-gap) of collective shear waves $k_{g}$ as a function of temperature in the subcritical liquid phase of Argon between $95$ K and $115$ K. The data are taken from Ref. Yang et al. (2017). For convenience, the data of the fraction of unstable INM $f_{u}$ are reproduced in the same panel. (b) The $k$-gap as a function of the fraction of unstable modes $f_{u}$ and as a function of the self-diffusion constant $D$.
The presence of unstable modes, $f_{u}\neq 0$, has been often identified as a hallmark of fluidity and a typical feature of liquids. As we have already observed, this is not that simple since a portion of unstable modes survive in the solid phase (at least near to melting) and a more refined classification of the unstable INM is needed to connect them directly with “fluidity” (see e.g. Bembenek and Laird (1995); Clapa et al. (2012a); Sciortino and Tartaglia (1997); La Nave et al. (2000, 2001); Chowdhary and Keyes (2002); Gezelter et al. (1997); Li and Keyes (1997) for the several criteria proposed). Nevertheless, we find it interesting to compare the fraction of unstable modes and the self-diffusion constant to another measure of “fluidity” known as the $k$-gap. As originally advocated by Maxwell and Frenkel Frenkel (1946), and later extensively re-discussed Trachenko and Brazhkin (2015b); Baggioli et al. (2020), collective shear waves in liquids do not display a linear dispersion relation, as in solids, due to the vanishing of the static shear modulus. On the other hand, the real part of their dispersion (i.e., their energy) follows an approximate square-root form $\mathrm{Re}(\omega)=v\sqrt{k^{2}-k_{g}^{2}}$ where $k_{g}$ is known in the literature as $k$-gap. This structure has been confirmed in several simulations (e.g., Yang et al. (2017); Khusnutdinoff et al. (2020); Ge et al. (2023)) and has been also observed in few experimental systems Nosenko et al. (2006); Jiang et al. (2024). Interestingly, the value of $k_{g}$ is rather universal when compared to the Wigner-Seitz radius Baggioli (2021). $k_{g}$ is expected to vanish at the melting transition, below which the collective shear waves reacquire their propagating nature as predicted by hydrodynamics Martin et al. (1972) and elasticity theory Chaikin and Lubensky (1995). This suggests that somehow $k_{g}$ should relate to the self-diffusion constant and the fraction of “diffusive” unstable INM that also vanish in the solid phase. Nevertheless, to the best of our knowledge, how this net of relations is concretely realized has never been investigated.
Taking advantage of the existing data for $k_{g}$ for subcritical liquid Ar Yang et al. (2017), we have attempted a raw comparison among these quantities. The results are shown in Fig.6. In panel a, we have presented the data for $k_{g}$ as a function of temperature in the liquid phase. As expected, $k_{g}$ grows monotonically with temperature compatible with the idea that $k_{g}\propto 1/\tau_{m}$, with $\tau_{m}$ being the Maxwell time. The latter decreases with temperature and make $k_{g}$ to increase with $T$. For comparison, in the same panel, we also show the fraction of unstable modes as a function of temperature, to highlight the very similar trend. In panel b of Fig. 6, we plot $k_{g}$ as a function of $D$ and $f_{u}$. In both cases, we observed a very strong linear correlation between these quantities. This is consistent with the idea that all of them are a measure of “fluidity”.
In order to find a more robust, and possibly universal relation between these three quantities, few problems need to be solved. First, we notice that despite $f_{u}$ is a dimensionless quantity, $k_{g}$ and $D$ are not. Therefore, any universal relation between $f_{u}$ and $k_{g}$ for example needs necessarily to involve another dimensionful scale. The same argument holds for a direct relation between $k_{g}$, which has units of $[1/L]$, and $D$, which has units of $[L^{2}/t]$. Moreover, as already anticipated, $k_{g},D$ are expected to vanish when the solid phase is approached, while $f_{u}$ does not. In general, it would be interesting to investigate further this point, considering also other more refined measures of $f_{u}$ that do vanish in the solid phase.
IV What can we learn from instantaneous normal modes
In this Section, we have selected a set of interesting topics and open points surrounding our initial question of what we can learn from INM.
IV.1 Fraction of unstable modes
The fraction of unstable modes defines the percentage of negative eigenvalues (or equivalently imaginary frequencies) in the INM spectrum and it is a function of all thermodynamic variables, such as temperature and pressure.
x
Figure 7: The fraction of unstable modes as a function of the temperature. The numbers correspond to the power-law $\alpha$ described in the main text. (a) For Ar, Xe, Ga, Pb the data are normalized by their melting temperature $T_{m}$ and the value of $f_{u}$ at such temperature. (b) For N2 and CS2 the data are normalized by the lowest temperature available, $T_{0}$.
In all liquids, the fraction of unstable modes increases with temperature. A larger kinetic energy corresponds to a larger probability of exploring regions of the potential landscape with locally negative curvature, such as saddle points or maxima, moving the dynamics away from the bottom of the local minima. Despite this might seem natural from a physical perspective, how $f_{u}$ increases with temperature is much less trivial. In order to provide valuable information regarding the temperature trend of $f_{u}$, we have collected all our results in Fig.7. In panel a, we show $f_{u}$ for Ar, Xe, Ga and Pb. For a better comparison, we normalized the data using the melting temperature of each of these liquids $T_{m}$ and the corresponding value of $f_{u}$ at melting. We observe that $f_{u}(T)$ is well described in all cases by a power-law function $f_{u}(T)\sim T^{\alpha}$, where nevertheless $\alpha$ depends on the specific liquid chosen. More precisely we find that $\alpha$ is $1.09,1.06,0.84,0.71$ for Ar, Xe, Ga and Pb respectively. For N2 and CS2 we do not have data up to the melting temperature and we therefore decided to normalize the data by the minimum temperature available, $T_{0}$, and show the corresponding results in a separate panel, panel b in Fig.7. There, we find again a trend consistent with a single power-law, where the power $\alpha$ is $1.14$ for N2 and $0.88$ for CS2.
It is noteworthy that the power-law $\alpha$ is close to $1$ for all the liquids considered. This is in good agreement with the predictions from the random energy model (see e.g. Keyes (2005)), according to which $f_{u}(T)=\frac{T}{\delta 2\pi}e^{-\delta^{2}/(4T^{2})}$ where $\delta$ is an $\mathcal{O}(1)$ parameter describing the width of the energy distribution. We have tried to fit the data with the expression from the random energy model with a single free parameter $\delta$ and found that it does not capture correctly the trend of the numerical data.
We notice that the exponent $\alpha$ measures the rate at which a given liquid develops unstable modes by increasing the temperature. This exponent is possibly related both to dynamical properties of the liquids but, more fundamentally, to the dynamics of its potential energy landscape. In this direction, it would be interesting to explore which properties of a liquid determines determine such exponent and whether the latter displays any strong correlation with any other measurable quantity.
IV.2 Low frequency linear scaling
As already anticipated, another universal feature of the INM density of states is its linear low-frequency behavior of the form
$\rho(\omega)=a(T)|\omega|+\dots$ (13)
This behavior can be derived using several theoretical arguments Xu and Stratt (1989); Wu and Loring (1993); Wan and Stratt (1994); Zaccone and Baggioli (2021); Schirmacher et al. (2022); Li and Keyes (1999); Zürcher and Keyes (1997); Wu and Loring (1992) and it is directly related to the presence of unstable modes with imaginary frequency. In all the systems considered in this manuscript, that include Ar, Xe, Ga, Pb, N2 and CS2, the INM DOS follow the linear form presented in Eq.(13). In general terms, much less is known about the temperature dependence of the parameter $a(T)$ that characterizes the slope of this linear trend.
In a series of works Keyes (1994b); Li and Keyes (1999) (see Keyes (2005) for a review), Keyes and collaborators proposed a closed-form formula for this parameter that, under some reasonable approximations, takes the following form
$a(T)\propto\left[f_{u}^{\text{max}}-f_{u}(T)\right]\,e^{-\langle E\rangle/k_{B}T}\,.$ (14)
Here, $f_{u}^{\text{max}}$ is the maximum value for the fraction of unstable modes, and depends on the details of the potential landscape. On the other hand, $\langle E\rangle$ is an average energy scale that is supposedly related to hopping processes and the anharmonic nature of the potential landscape. As seen in the previous Section, the fraction of unstable modes grows with temperature. This implies that the first term in Eq.(14) is a term that decreases with temperature. The other contribution to Eq.(14) is an exponential growth of the scaling upon increasing the temperature, whose growth rate is controlled by the energy scale $\langle E\rangle$. Because of the competition between these two terms, in general Eq.(14) predicts that $a(T)$ grows/decreases with temperature depending on which of the terms therein dominates. These two behaviors have been both observed in simulations. For a Lennard Jones liquid Keyes (1994b), the energy scale $\langle E\rangle$ is very low; hence, the first term in Eq.(14) dominates the behavior of $a(T)$ that decreases with temperature. The same is observed for bulk polymers Zhang et al. (2019). On the contrary, for many other systems, including CS2 Li and Keyes (1999) and water Jin et al. (2024), $a(T)$ grows with temperature following a form compatible with the exponential term in Eq.(14).
Figure 8: The INM slope extracted from the experimental data using Eq.(13) as a function of the inverse temperature $1/T$. The red dashed line in panels (a)-(d) is the result of an exponential fit $a(t)\propto\exp(-\langle E\rangle/k_{B}T)$. For all systems, the data points are above the melting temperature.
We have analyzed the slope parameter $a(T)$ for our six different liquid systems. For Ar, Xe, N2 and CS2 we observe that the slope $a(T)$ clearly increases with temperature. Moreover, for these systems the temperature dependence of the slope is compatible with an Arrhenius-like exponential form $a(t)\propto\exp(-\langle E\rangle/k_{B}T)$, where the activation energy is given by:
[…]
We notice that the reported value of $\langle E\rangle\approx 13\,\text{meV}$ for water Jin et al. (2024) is of the same order of magnitude of the values found above. As we will see in the next Section, these activation energies are consistently smaller than those governing the Arrhenius behavior of the corresponding diffusion constants. For example, for $N_{2}$ the activation energy characterizing the slope $a(T)$ is about $20$ times smaller than the one describing $D(T)$. This implies that, despite the same functional form, it is very unlikely that the microscopic origin of it is the same for the slope $a(T)$ and the diffusion constant $D(T)$. Nevertheless, we notice that all the normal modes analysis, including therefore the determination of $a(T)$, is performed using the harmonic approximation. On the contrary, the determination of $D(T)$ does not rely on this approximation. It is therefore conceivable (similar arguments have been made in Jin et al. (2024) by comparing the INM slope and the experimental slope) that the strong liquid anharmonicities are the responsible for this difference.
Finally, we clarify that, to the best of our knowledge, no existing theoretical framework to determine from first principles $\langle E\rangle$ exists. Our numerical results could be useful in this direction to achieve such a theoretical model and directly test it with simulations.
On the other hand, as shown in panels (e)-(f) in Fig.8, the temperature dependence of the slope in Ga and Pb shows a completely different trend. In particular, within the accuracy of our numerics, the slope seems rather independent of temperature and shows an approximate constant behavior. At the moment, we do not have any explanation for this behavior but we can already offer two insights. (I) The numerical data are certainly incompatible with an exponential behavior $a(t)\propto\exp(-\langle E\rangle/k_{B}T)$ as found for the other simpler liquids. (II) Both Ga and Pb have in common the fact that they are liquid metals, providing a possible hint to rationalize this different trend.
IV.3 Self-diffusion constant
One of the initial motivations to introduce and study instantaneous normal modes was the idea of reconstructing the liquid self-diffusion constant in terms of liquid normal modes Zwanzig (1983). Keyes and collaborators have proposed that the self-diffusion constant in liquids could be related to the unstable instantaneous normal modes Madan et al. (1990b), and that a somehow simple relation between $D$ and $f_{u}$ could exist,
$D(T)=\text{const.}\,\langle\omega_{u}\rangle f_{u}.$ (16)
Here, $\langle\omega_{u}\rangle$ is the average unstable frequency. We notice that in Eq.(16), “const.” is not a dimensionless number but it is assumed to be temperature independent.
Despite INM describes diffusion in liquids well Keyes (1997b, 1994b); Madan et al. (1990b); Keyes (1995), several open questions remain. First, it is by now well-known that the total fraction of unstable modes cannot be a direct source for self-diffusion since $D=0$ in solids but $f_{u}$ is not. As already mentioned above, this has motivated a large number of studies to find out which subset of $f_{u}$ is actually responsible for self-diffusion (e.g., La Nave et al. (2001); Sciortino and Tartaglia (1997) and discussions therein). Moreover, in several instances (e.g., CS2) it has been shown that Eq.(16) does not hold and a “stretched” version of Eq.(16) has to be considered (see e.g. Li and Keyes (1999)),
$D(T)\sim\left(\langle\omega_{u}\rangle f_{u}\right)^{\gamma}.$ (17)
We have computed the temperature dependent self-diffusion constant $D(T)$ for our six liquid systems. The behavior of $D$ as a function of temperature is presented in Arrhenius form in Appendix 4. All data are consistent with an Arrhenius behavior,
$D(T)\propto e^{-\mathcal{E}/T}.$ (18)
Our numerical fits give:
[…]
Figure 9: The self-diffusion constant as a function of the fraction of unstable modes $f_{u}$. The numbers correspond to the power-law $\beta$ described in the main text. (a) For Ar, Xe, Ga, Pb the data are normalized by their melting temperature $T_{m}$ and the value of $D$ at such temperature. (b) For N2 and CS2 the data are normalized by the lowest temperature available, $T_{0}$.
As already emphasized in the previous Section, the activation energy $\mathcal{E}$ governing the Arrhenius behavior in $D(T)$ is consistently larger than the one describing the increase of the INM slope in the corresponding systems. Not only that, but even the trend of the values between the different liquids is not the same. For example, $\langle E\rangle_{\text{Ar}}>\langle E\rangle_{\text{CS}_{2}}$ but $\mathcal{E}_{\text{Ar}}<\mathcal{E}_{\text{CS}_{2}}$. This implies that either the two energy scales reflect very different properties of a liquid or anharmonicities, that are neglected in the determination of $\langle E\rangle$, play a fundamental role here and are the origin of this discrepancy.
In order to investigate further the relation between the self-diffusion constant $D$ and the fraction of unstable modes, in Fig.9 we have plotted $D$ as a function of $f_{u}$. In order to compare the different systems, we have normalized both quantities by subtracting their value at the melting temperature $T_{m}$ (for Ar, Xe, Ga and Pb) or at the minimum temperature available $T_{0}$ (for N2 and CS2). For all the systems considered, we observe a power law scaling, with the power $\beta$ ranging from $1.15$ in Xe to $1.96$ in CS2. We notice that most of the powers obtained by fitting are not far from the value of $1$, that would be consistent with Eq.(16) and the analysis in Li and Keyes (1999). Nevertheless, we also observe important deviations and we do not find any universal scaling.
IV.4 Low temperature solid-like limit
Let us consider a liquid. Its INM DOS is linear in the frequency. On the other hand, Debye taught us that for a zero temperature ideal crystal, the corresponding vibrational density of states has a quadratic behavior, known as Debye’s law Kittel (2021). The contrast between these two opposite trends has been observed experimentally in selenium Phillips et al. (1989), heavy water Dawidowski et al. (2000) and water Jin et al. (2024). Less is known about how the INM DOS interpolates between these two behaviors by lowering the temperature and approaching the ideal crystalline phase. As already mentioned, the presence of unstable modes (no matter how few) implies that a linear scaling appears in the INM DOS. Therefore, as soon as unstable mode are present in the spectrum and $f_{u}\neq 0$, the INM DOS will be linear at low enough frequencies. By decreasing temperature, the fraction of unstable modes diminishes and the linear scaling is expected to get confined to very low frequency and eventually give room to the Debye scaling, that is allowed to appear only when $f_{u}=0$.
Figure 10: (a) The experimental density of states for liquid gallium at $500$ K reproduced with permission from Ref.Stamper et al. (2022). (b) The INM DOS of Gallium simulated at different temperatures. The first-order melting temperature is between 320 K and 340 K. (c) The density of state function $S(\omega)$ derived from the Fourier transform of the VACF for the same simulated Gallium system.
In order to investigate this point, we have chosen Gallium as a benchmark example. As a reference, in Fig.10, we present the experimental data for the density of states at $500$ K from Ref.Stamper et al. (2022) together with our simulation results for the INM density of states $\rho(\omega)$ and the density of states function $S(\omega)$. Among the various features emerging from this comparison, we notice that at low temperature (e.g., $T=300$ K), $S(\omega)$ shows already a clear Debye scaling at low frequency and a vanishing diffusion constant ($S(0)=0$). On the contrary, the stable branche of the INM DOS is still strongly linear at low frequency, consistent with a large number of unstable mode still present at such a temperature.
Figure 11: Simulated Gallium at $1$ K. (a) The comparison between the stable branch of the INM DOS $\rho(\omega)$ and the density of states function $S(\omega)$ derived from VACF. (b) Zoom of the low frequency region of the stable branch of the INM DOS. Red and blue dashed lines guide the eyes towards the Debye scaling $\omega^{2}$ (blue) and the linear scaling (red). The corresponding vertical dotted lines mark the regions where these scalings appear.
In order to test the emergence of the Debye scaling in the INM spectrum, we have lowered the temperature to a value of $T=1$ K. The results of this numerical analysis for $\rho(\omega)$ and $S(\omega)$ are presented in Fig.11. The density of state function $S(\omega)$ presents a Debye scaling $S(\omega)\propto\omega^{2}$ up to approximately $8$ meV and it exhibits two distinct peaks at $\approx 10$ meV and $18$ meV. Moving to the INM analysis, at $T=1$ K, we observe that the number of unstable modes in the INM spectrum becomes very small. The shape of $\rho(\omega)$ is now very consistent with the density of state function $S(\omega)$, as shown in panel (a) of Fig.11. The INM analysis is able to capture well the position of the two aforementioned peaks, despite their width is slightly larger and intensity smaller. This seemingly good agreement between the two quantities is a consequence of the fact that the harmonic approximation becomes more and more reliable at low temperature. In fact, at exactly $T=0$, the two computations should give exactly the same results, up to numerical precision.
In order to check this point further, in panel (b) of Fig.11, we have zoomed the stable part of the INM spectrum of Gallium at $T=1$ K and presented it in a log-log scale to highlight possible power-law scalings. The data present a clear linear scaling up to approximately $4.5$ meV, that is a consequence of the remaining unstable modes (and anharmonicities). On the other hand, around $\approx 4.5$ meV we observe a clear crossover to a larger power that is consistent with the Debye scaling $\rho(\omega)\propto\omega^{2}$ and that persists to $\approx 8$ meV. This reveals the approach of the solid-like behavior to low frequency by decreasing the temperature. By decreasing further $T$, one expects that the crossover frequency will move towards the origin, and finally vanish at $T=0$ when $f_{u}=0$. We notice that an alternative method to observed this trend, and the emergence of the Debye scaling, consists in nano-confining the liquid, as experimentally performed in Yu et al. (2024). We also notice that the presence of unstable modes in finite temperature (amorphous) solids has been verified experimentally, see for example Vaibhav et al. (2024).
IV.5 Zero eigenvalue cusp singularity
Following previous observations Sastry et al. (2001); Taraskin and Elliott (2002), Refs. Schirmacher et al. (2022); Mossa et al. (2023) have recently pointed out that the frequency representation of the INM density of states $\rho(\omega)$ might hide important information that are contained in the corresponding eigenvalue distribution $\upsilon(\lambda)$. Their arguments are supported by numerical simulations in a soft-sphere liquid model (data reproduced in panel (a) of Fig.14). Since $\lambda=\omega^{2}$, the eigenvalue distribution and the INM DOS are simply related via the Jacobian,
$\rho(\omega)=2|\omega|\upsilon(\lambda=\omega^{2}).$ (20)
Mathematically, it appears therefore unreasonable that one of the two quantities contain more information than the other. Nevertheless, in this Section we entertain this possibility and analyze in detail the arguments of Schirmacher et al. (2022); Mossa et al. (2023) in a large set of liquid models. First, we would like to point out that the statement that the INM density of states is universally linear at low frequencies is equivalent to the statement that $\rho(0)\neq 0$, that is certainly true. In other words, the presence of a cusp at $\lambda=0$ is not incompatible with a low-frequency form $\rho(\omega)=a(T)\omega+\dots$, observed in many instances.
Here, we have obtained the eigenvalue distribution $\upsilon(\lambda)$ for six different liquids (Ar, Xe, Pb, Ga, N2 and CS2). The results are shown in a large range of temperatures in Fig.12. The cusp-like singularity at $\lambda=0$ is present in all the systems considered. The temperature dependence of the value of $\upsilon(\lambda)$ at the cusp, $\lambda=0$ does not bring any new information since $\upsilon(0)=a(T)/2$, where $a(T)$ is the slope of the low-frequency regime in the frequency representation $\rho(\omega)$. The asymmetry character of the cusp-like singularity is more interesting and relates to the different behavior of the stable and unstable INMs upon moving temperature.
In order to characterize these features in a more quantitative way, we resort to the empirical function proposed in Mossa et al. (2023),
$\upsilon(\lambda)=\frac{e^{-a\lambda}(c_{0}+c_{1}\lambda)+e^{a\lambda}(c_{2}+c_{3}\lambda)}{e^{-a\lambda}+e^{a\lambda}}$ (21)
where $c_{1},c_{3}$ are referred to as “slopes” of the cusp-singularity. More precisely, $c_{1}$ is the slope of the stable branch and $c_{3}$ that of the unstable one.
Figure 12: The eigenvalue distribution $\upsilon(\lambda)$ for different temperatures. From top left to bottom right: Ar (40 bar), Xe ($1.01$ bar), Pb($1.01$ bar), Ga($1.01$ bar), N2 ($70.92$ bar) and CS2 ($1.01$ bar).
Figure 13: (a) The temperature dependence of the “slopes” of the eigenvalue distribution $\upsilon(\lambda)$ for CS2 extracted using the empirical formula Eq.(21). (b) The slope difference as a function of temperature. The dashes line guides the eyes towards the linear scaling observed.
In panel (a) of Fig.13, we show the temperature behavior of the fitted stable (red) and unstable (black) slopes as a function of temperature for the case of CS2. The slopes are extracted from the numerical data using Eq.(21). Similar results are obtained for the different liquids shown in Fig.12. We observe that the positive slope of the unstable branch grows with temperature while the negative one, corresponding to the unstable part of the spectrum, moves towards larger negative values by increasing $T$. This trend is opposite of what reported in Schirmacher et al. (2022); Mossa et al. (2023) for the soft sphere model. The reason for this difference is ultimately the opposite behavior of $\upsilon(0)$ with temperature in the soft sphere model. In particular, as evident from Fig.1 in Schirmacher et al. (2022), $\upsilon(0)$ in the soft sphere model decreases by increasing temperature. In our language, this corresponds to an INM slope $a(T)$ decreasing with temperature, as for LJ liquids, but opposite to our more realistic liquid systems.
This contrast is then reflected in the temperature dependence of the difference of the slopes, that is presented in panel (b) of Fig.13. There, we observe that $\Delta_{\text{slope}}$ increases with temperature, following a trend approximately consistent with a linear power-law scaling. This is again very different from the $T^{-2/3}$ behavior observed in the soft sphere model Schirmacher et al. (2022); Mossa et al. (2023), not only for the precise power involved in the scaling, but even in the qualitative trend as well. Our findings imply that the $T^{-2/3}$ law found in Schirmacher et al. (2022); Mossa et al. (2023) is not universal and its value and significance should be revisited.
Finally, we would like emphasize that the theoretical model presented in Schirmacher et al. (2022); Mossa et al. (2023) is not the only one able to predict the presence of a cusp-like singularity in the eigenvalue distribution of liquids. Indeed, a much older model known as the “soft potential model of liquids” (SPM) Zürcher and Keyes (1997); Xu et al. (2023) already captures such a feature. We provide a direct proof of this statement in panel b of Fig.14. The presence of a cusp singularity at $\lambda=0$ is evident. For completeness, in the inset we also show the INM DOS that shows the aforementioned linear in frequency regime despite the cusp.
Figure 14: (a) The eigenvalue distribution $\upsilon(\lambda)$ for the soft-sphere liquid model presented in Mossa et al. (2023) and the corresponding fraction of unstable mode $f_{u}$ as a function of temperature in reduced LJ units. $T_{g}$ is the glass temperature of the corresponding system. (b) The eigenvalue distribution $\upsilon(\lambda)$ obtained from the theoretical soft potential model of liquids Zürcher and Keyes (1997); Xu et al. (2023) by converting the calculated $\rho(\omega)$ to $\upsilon(\lambda)$ with the Jacobian relation. The inset shows the original INM DOS $\rho(\omega)$ with the dashed black lines indicating the universal linear regime $\rho(\omega)\propto\omega$ which emerges clearly despite the presence of the cusp-like singularity in $\upsilon(\lambda)$.
In this direction, it would be interesting to understand further the connection between the theory presented in Schirmacher et al. (2022); Mossa et al. (2023) and the older SPM Zürcher and Keyes (1997); Xu et al. (2023). We notice that the two frameworks start from completely different assumptions. The heterogeneous elasticity theory of Schirmacher et al. (2022); Mossa et al. (2023) follows a harmonic description, whose validity is arguable in the liquid phase, especially at large temperatures. On the other hand, the SPM starts from a anharmonic picture of the dynamics that involves additional low-energy modes. Also in this case, the nature and origin of these modes in liquids remains an open question. String-like excitations Donati et al. (1998) can represent a candidate for these anharmonic excitations.
V Conclusions
What do instantaneous normal modes tell us about liquid dynamics? Our work revolves around this question.
We have considered six different liquids, including Ar, Xe, Pb, Ga, N2 and CS2, in order to provide a (as much as possible) complete survey of INM properties and their relation to other standard probe for liquid dynamics, such as the velocity auto-correlation function, the self-diffusion constant and the dispersion of collective low-energy modes.
The highlights of our extensive analysis are collected below.
  • The linear scaling of the low-frequency INM DOS has been confirmed for all six liquids. This validates the hypothesis that, as far as the fraction of unstable mode is finite, the Debye scaling does not emerge in the INM spectrum, not even in the finite temperature solid phase. Moreover, this linear scaling has been proven to coexist with a cusp-like singularity in the eigenvalue spectrum, showing that, contrary to recent claims in the literature Schirmacher et al. (2022), the two features are not incompatible.
  • The slope of the INM spectrum in the linear scaling regime $a(T)$ grows with temperature in simple liquids, following an Arrhenius-like exponential form, as previously demonstrated in Jin et al. (2024). On the contrary, metallic liquids display an approximate constant slope, that results (at least in the range of $T$ investigated) independent of temperature.
  • Both the fraction of unstable modes, but also the slope of the INM spectrum in the linear scaling regime $a(T)$, display a sharp jump at the first-order melting transition but no signature at the Frenkel dynamical crossover.
  • The fraction of unstable modes, the momentum gap of shear collective waves (k-gap) and the self-diffusion constant display a very strong correlation in the whole liquid phase, confirming their role as “probes of fluidity”.
  • The temperature dependence of the fraction of unstable modes, but also the relation between the self-diffusion constant and the latter, do not appear to be universal in the different liquid systems considered.
  • Against common expectations, crystals at finite temperature display a rather large number of unstable modes, that becomes negligible only at very low temperature. As a consequence, the scaling of the INM spectrum remains linear. At low temperature, a crossover scaling between a solid-like Debye scaling and a linear regime emerges in the INM spectrum, where the crossover frequency scale decreases with $T$.
  • In all liquids considered, the eigenvalue distribution displays a cusp-like singularity at $\lambda=0$. This feature can be explained both using heterogeneous-elasticity theory Schirmacher et al. (2022); Mossa et al. (2023), but also using the soft potential model of liquids Zürcher and Keyes (1997); Xu et al. (2023). Moreover, the temperature dependence of the intermediate slopes and their difference as a function of temperature do not show the behavior previously reported in the literature for a simulated soft sphere liquid model Schirmacher et al. (2022); Mossa et al. (2023).
We conclude by noticing that many of these observations are yet not understood nor rationalized from a theoretical perspective, leaving an important hole in our understanding of liquid dynamics and interesting research directions for the near future. We hope that our work will play a motivating role in this direction.
Acknowledgements.
We would like to thank T. Keyes, J. Douglas, A. Zaccone, Y. Wang, J. Zhang, D. Yu, C. Stamper, J. Moon, Y. Yu, L. Hong and H. Xu for many discussions about INM and liquids. We also would like to thank D. K. Belashchenko for instructions about the Gallium potential. We also thank T. Keyes, J. Douglas and J. Moon for useful comments on a preliminary version of this manuscript. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. X. F. would like to thank the support provided by the National Natural Science Foundation of China (No. 52201016) and acknowledge the support of the State Key Laboratory of Clean and Efficient Turbomachinery Power Equipment for providing HPC resources.
Appendix 1 Additional data for Argon
In the main text, we have shown the data for Argon only along the constant pressure line with $P=40$ bar. In Fig. A1, we report additional data at other pressures that confirm the findings discussed in the main text.
Figure A1: Panels (a) and (b) show the INM DOS $\rho(\omega)$ and the density of state function $S(\omega)$ for Argon at $1000$ bar as a function of temperature. Panels (c) and (d) show the same quantities at fixed temperature, $T=120$ K, as a function of the pressure.
Appendix 2 Xenon, Nitrogen and more liquid systems
Argon has been taken in the main text as a benchmark example for our analysis. Here, we report additional data for the other liquids discussed in this manuscript. Fig. A2 reports additional data for the INM DOS at different temperatures for Xe, Pb, N2 and CS2. Fig. A3 shows the density of state function at different temperature for Xe and Pb.
Figure A2: INM DOS $\rho(\omega)$ as a function of temperature for Xe (a), Pb (b), N2 (c) and CS2 (d).
Figure A3: Density of state function $S(\omega)$ for Xe (a) and Pb (b) at different temperatures.
Appendix 3 Self-diffusion constant
In Fig. A4 we show the Arrhenius plot for the self-diffusion constant of the six different liquid systems studied in our work. The corresponding activation energies are reported Eq. (19) in the main text.
Figure A4: Arrhenius plot for the self-diffusion constant $D$ as a function of temperature. Different panels correspond to the six different liquids considered in this work. The red dashed lines indicate the Arrhenius fit.
Appendix 4 Slopes of the eigenvalue distribution $\nu(\lambda)$
In Fig. A5, we show the temperature behavior of the fitted stable and unstable slopes discussed in the main text for the other systems, including Ar, Xe, Pb and Ga.
Figure A5: The “slopes” of the eigenvalue distribution $\upsilon(\lambda)$ as a function of temperature for Ar (a), Xe (b), Pb (c) and Ga (d).